# Causal Concept Embedding Models: Beyond Causal Opacity in Deep Learning

### v1

@@ Abstract
Causal opacity denotes the difficulty in understanding the “hidden” causal structure underlying a deep neural network’s (DNN) reasoning. This leads to the inability to rely on and verify state-of-the-art DNN-based systems especially in high-stakes scenarios. For this reason, causal opacity represents a key open challenge at the intersection of deep learning, interpretability, and causality. This work addresses this gap by introducing Causal Concept Embedding Models (Causal CEMs), a class of interpretable models whose decision-making process is causally transparent by design. The results of our experiments show that Causal CEMs can: (i) match the generalization performance of causally-opaque models, (ii) support the analysis of interventional and counterfactual scenarios, thereby improving the model’s causal interpretability and supporting the effective verification of its reliability and fairness, and (iii) enable human-in-the-loop corrections to mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also accuracy of the explanation provided for a specific instance.
@@ 1 Introduction
Deep Learning (DL) models have a pervasive impact on many areas of contemporary research and society [1]. Despite this success, there is growing concern about the widespread real-world application of DL, particularly in sensitive domains [2]. These concerns are partly due to the lack of causal explainability of these models, which undermine their robustness, fairness and generalisability [3]. Causal explainability, in particular, is a multi-faced issue, which includes a variety of different, albeit related, problems [4]. One such problem is that of causal discovery and concerns the possibility of using a model to detect and understand the causal mechanisms of the data generating process [5, 6, 3, 7] (see Figure 1(a)). An equally important problem is that of causal opacity, which denotes the difficulty of users to grasp and understand the “hidden” causal structure characterising a model’s inference and behaviour (see Figure 1(b)).
Causal opacity can be better understood in the light of Pearl’s framework of causality [8, 6], which measures an agent’s causal understanding in terms of their ability to answer what-if type of questions, and specifically interventional and counterfactual questions. In this regard, causal opacity depends on the ability of users to answer interventional and counterfactual questions regarding the structure of a model’s inference (e.g., “what happens if I fix the feature age to a value greater than $50$?”). This capacity is vital for assessing a model’s reliability and robustness and plays a central role in establishing its fairness, especially within the popular framework of counterfactual fairness  [9]. For instance, in a loan recommendation system, ensuring fairness might involve verifying that sensitive demographic attributes, like gender or ethnicity, have no causal influence on model’s decision-making process. This task can be solved by checking that such predictions are robust (i.e., do not vary) under interventional and counterfactual adjustments. This ultimately requires answering interventional and counterfactual questions such as “will the decision change if I vary the applicant’s ethnicity?”.
$x$Raw Features $v_{o}$arget Data Generation (Causal Discovery) Model Inference (Causal Opacity)
(a) Causal explainability
@@
(b) Black Box
@@
(d) Causal CEMs
@@
Figure 1: (a) Causal explainability has two distinct dimensions: causal discovery, where we aim to identify and understand the causal mechanisms underlying a data-generating process, and causal opacity, where we aim to understand the causal structure of a model’s inference to verify its robustness, generalisability, and fairness. (b) Standard DL models are black boxes in the sense that the causal structure of their mapping from raw input features (e.g., pixels of an image) to the target remains opaque. (c) In Concept Bottleneck Models (CBM), high-level human-interpretable concepts are first extracted through an encoder $g$ and then used to predict the target. Although CBMs are semantically interpretable, the causal structure of the model’s inference assumes a straightforward causal structure where concepts are causally independent and are all direct causes of the target. (d) In Causal Concept Embedding Models (Causal CEMs), both the concepts’ semantics and the inference’s causal structure are transparent and interpretable.
To address causal opacity, the field of eXplainable AI (XAI) has developed a variety of techniques aimed at explaining the causal mechanisms underlying a model’s predictions [10, 11, 12]. Among the most common methods are feature attribution techniques, where an explanation aims to measure the causal relevance of each feature (or latent factor) for a model’s outcome [13, 14, 15, 16, 17, 18, 19, 20]. Although these techniques are promising, their focus on “raw” input features limits their applicability, especially on unstructured data (e.g., images) whose input features denote low-level attributes (e.g., pixels) lacking an understandable and contextually relevant meaning. For this reason, many works stress that good causal explanations should be concept-based [21], i.e., they should rely on the “high-level” features, or “concepts”, that a model infers from raw data when making predictions [3, 22]. State-of-the-art concept-based models—such as Concept Bottleneck Models [23] and concept-based counterfactuals [24, 25]—partially address this issue while being prone to a key limitation : they only capture direct counterfactual dependence [26], a weak form of causal dependence that assumes all concepts to be causally mutually independent and directly related to the target prediction (see Figure 1(c)). This oversimplifies an inference’s causal structure, which typically encompasses a complex, dense network of causal dependencies among various concepts. As a result, the problem of causal opacity still represents a key open challenge at the intersection of DL, causality, and XAI that currently limits the explainability, reliability, and verifiability of modern DL systems.
@@
To bridge this gap, we introduce Causal Concept Embedding Models (Causal CEMs, see Figure 1(d)), an interpretable concept-based architecture delivering inferences whose causal structure is transparent by design. The results of our experiments show that Causal CEMs can: (i) match the generalisation performance of state-of-the-art causally-opaque models, (ii) support the analysis of interventional and counterfactual scenarios, thereby improving the model’s causal interpretability and supporting the effective verification of its reliability and fairness, and (iii) enable human-in-the-loop corrections of mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also accuracy of the explanation provided for a specific instance.
@@ 2 Background
A causal model (CM) is a mathematical representation of the mechanisms (rules, laws) that explain how different variables of a given target-system causally influence each other. In computer science and statistics, the standard framework for causal modelling is that of Structural Causal Models (SCM) proposed by Pearl [8]. Formally, an SCM $M$ is a triplet $(\mathcal{U},\mathcal{V},\mathcal{F})$ where: $\mathcal{U}$ is a set of exogenous variables representing latent factors with a causal influence on the modelled target-system; $\mathcal{V}$ is a set of endogenous variables representing observable and measurable variables; $\mathcal{F}$ is a set of functions (or structural equations) describing the causal mechanisms, which determine the values of each endogenous variable $v_{i}\in\mathcal{V}$ by computing $v_{i}=f_{i}(u_{i},\text{pa}(v_{i}))$, where $u_{i}\in\mathcal{U}$ and $\text{pa}(v_{i})$ are, respectively, the set of exogenous and endogenous variables causally affecting $v_{i}\in\mathcal{V}$. Every SCM can be associated with a directed acyclic graph (DAG) whose nodes represent variables and edges represent direct causal connections. Interventions can be modelled in SCMs through the do-operator [27, 8] . This operator induces a modification on the model’s structure by changing the value of an endogenous variable which is fixed to a value $\kappa\in\mathbb{R}$ by external means, effectively breaking its original causal dependencies within the model. Formally, applying $do(v_{i}=\kappa)$ to a model $\mathcal{M}$ results in a new model $\mathcal{M}_{v_{i}=\kappa}$ which is identical to $\mathcal{M}$ except that the equation for $v_{i}$ is replaced with a constant value $\kappa$, removing all arrows into $v_{i}$ in the model annexed DAG.
@@
Concept-based models [28, 29, 30, 31, 32, 33, 34] are interpretable architectures that explain their predictions using high-level units of information (i.e., “concepts”). Most of these approaches can be formulated as a Concept Bottleneck Model (CBM) [23], an architecture where predictions are made by composing (i) a concept encoder $g:X\to C$ that maps samples $\mathbf{x}\in X\subseteq\mathbb{R}^{d}$ (e.g., pixels) to a set of $r$ concepts $c\in C\subseteq\{0,1\}^{r}$ (e.g., “red”, “round”), and (ii) a task predictor $f:C\to Y$ that maps predicted concepts to a set of $l$ tasks $y\in Y\subseteq\{0,1\}^{l}$ (e.g., labels “apple” or “tomato”). Each component $g_{i}$ and $f_{j}$ denotes the truth degree of the $i$-th concept and $j$-th task, respectively. Usually, concept-based models represent a concept $c_{i}$ using its predicted truth degree $\hat{c}_{i}\in[0,1]$. This representation, however, might significantly degrade task accuracy when the provided concepts are incomplete [35, 36]. To overcome this issue, Concept Embedding Models (CEMs) [36] use high-dimensional embeddings $\mathbf{\hat{c}}_{i}\in\mathbb{R}^{z}$ to represent concepts alongside their truth degrees $\hat{c}_{i}\in[0,1]$.
3 Causal Concept Embedding Models
To address the causal opacity of DL systems, we need a model that can answer both interventional and counterfactual causal queries. In order to enable human verification and control, the causal structure of the model’s inference should be based on high-level interpretable concepts (e.g., colours, shapes) as low-level attributes do not provide a controllable semantics (a specific pixel lacks meaningful semantics when it comes to causal queries)  [21]. In the absence of causal priors, we also need the model to learn causal dependencies among high-level features from available data without making strong assumptions such as direct counterfactual dependence [26]. Finally, the model should not sacrifice predictive generalisation performance in exchange for causal interpretability. To this end, we introduce Causal Concept Embedding Models (Causal CEMs, Figure 1(d)), a class of concept-based architectures delivering causally transparent and robust inferences. In this section, we present the Causal CEMs blueprint (Section 3.1) and training process (Section 3.2).
@@ 3.1 Blueprint
In the absence of causal priors, Causal CEMs need to learn causal dependencies between high-level features that lead to the model becoming more effective in solving its designated task. As a result, we can formally describe Causal CEMs using a generalised probabilistic graphical model (PGM) that extends a traditional PGM by allowing cycles:
Definition 3.1 (Causal Concept Embedding Model).
Given an observed input feature $x$, a set of $k\in\mathbb{N}$ latent factors $u_{i}\in\mathcal{U}$ each associated with a high level interpretable variable $v_{i}\in\mathcal{V}$, a Causal Concept Embedding Model is the generalised probabilistic graphical model (PGM) $G=(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}=\{x\}\cup\mathcal{U}\cup\mathcal{V}$ and edges =$\mathcal{E}=\big{\{}(x,u_{i})\mid u_{i}\in\mathcal{U}\big{\}}\cup\big{\{}(u_{i},v_{i})\mid i\in\{1,\cdots,k\}\big{\}}\cup\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$ which represents the joint conditional distribution $p(v,u\mid x)$:
@@
The cyclical nature of Causal CEMs comes from the necessity to model all possible dependencies among variables $v_{i}$. Notice that the model is uniquely identified by the set of all conditional probability distributions corresponding to the arrows in the graph. Unfortunately, in generalised PGMs, the model does not easily factorise in terms of such distributions due to the cycles. To deal with cycles while maintaining the independencies induced by the graph structure, we can use an unfolding semantics for cyclical PGMs [37]. This semantics is based on the choice of a “cutset” i.e., a specific set of nodes $\mathcal{Q}\subseteq\mathcal{N}$ in the PGM such that every cycle in the PGM contains at least one node in $\mathcal{Q}$. Intuitively, by unfolding the nodes in the cutset, all cycles are broken leaving us with a standard acyclical PGM. The consistency between the semantics of the original cyclical PGM and the unfolded acyclical PGM is only valid in the limit of infinite unfolding [37]. However, when computing the likelihood of an observed complete set of variables $v_{i}\in\mathcal{V}$, modelling one single unfolding (i.e., a single transition) suffices for learning the conditional probability distributions among the variables in $\mathcal{V}$, as all the variables become conditionally independent on each other. As a result, we can define a Dissected Causal CEM as the one-step unfolding of the Causal CEM in Definition 3.1:
Definition 3.2 (Dissected Causal CEM).
Given a Causal CEM, let $\mathcal{V}^{\prime}=\mathcal{V}$ be the cutset. Then, the dissected Causal CEM $\mathcal{G}=(\mathcal{N}\cup\mathcal{V}^{\prime},\mathcal{E}_{\mathcal{V}^{\prime}})$ is an acyclic PGM obtained by extending the generalised PGM by (i) adding a copy of all cutset nodes $\mathcal{V}^{\prime}=\{v_{i}\mid v_{i}\in\mathcal{V}\}$, (ii) adding a new set of edges directed from parents of cutset’s nodes to the generated copies $\mathcal{V}^{\prime}$ i.e., $\mathcal{E}_{\mathcal{V}^{\prime}}=\{(a,b)\mid(a,b)\in\mathcal{E},b\in\mathcal{V}^{\prime}\}\cup\{(a,b)\mid(a,b)\in\mathcal{E},b\notin\mathcal{V}^{\prime}\}$, and (iii) defining an initial probability distribution for the new copies $p(v^{\prime}|u)$ given the latent variables. The resulting PGM, factorised as $p(v,v^{\prime},u\mid x)=p(v\mid v^{\prime},u)p(v^{\prime}\mid u)p(u\mid x)$, is:

### v4

# Causal Concept Graph Models: Beyond Causal Opacity in Deep Learning

@@ Abstract
Causal opacity denotes the difficulty in understanding the “hidden” causal structure underlying the decisions of deep neural network (DNN) models. This leads to the inability to rely on and verify state-of-the-art DNN-based systems, especially in high-stakes scenarios. For this reason, circumventing causal opacity in DNNs represents a key open challenge at the intersection of deep learning, interpretability, and causality. This work addresses this gap by introducing Causal Concept Graph Models (Causal CGMs), a class of interpretable models whose decision-making process is causally transparent by design. Our experiments show that Causal CGMs can: (i) match the generalisation performance of causally opaque models, (ii) enable human-in-the-loop corrections to mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also the reliability of the explanations provided for specific instances, and (iii) support the analysis of interventional and counterfactual scenarios, thereby improving the model’s causal interpretability and supporting the effective verification of its reliability and fairness.
@@ 1 Introduction
Deep Learning (DL) models (Fig. 1(a)) are causally opaque when applied to unstructured data like images, as interventions on low-level features, such as pixels, are not meaningful (Rudin, 2019; Ghorbani et al., 2019) and therefore do not enable us to easily answer any of these sorts of questions. For example, asking “what happens if I fix the intensity of the pixel $x_{i}$ to $k$?” does little to reveal whether a DNN is exploiting protected attributes like “gender” or “race” when classifying an image of a person as a “doctor”. Yet, understanding when these sorts of high-level attributes are misused to make predictions is crucial for the proper verification and vetting of models (Kusner et al., 2017), particularly in high-impact real-world tasks (e.g., medical admission recommendations).
To address this issue, Concept Bottleneck Models (CBMs) (Koh et al., 2020) have been designed to support interventions on high-level interpretable features known as “concepts” (Kim et al., 2018). These concepts, whose semantics are aligned with units of information experts would use to solve the same task, allow for more intuitive interventions and counterfactual analysis (Wachter et al., 2017; Abid et al., 2021). This is because CBMs shift from reasoning based on raw features to explicitly operating on interpretable variables such as “age”, “income”, “gender”, or “BMI”.
@@
(b) Standard CBMs
[…]
(c) Causal CGMs
@@
Figure 1: (a) Standard DL models are black boxes in the sense that the causal structure of their mapping from raw input features (e.g., pixels of an image) to the target remains opaque. (b) In Concept Bottleneck Models (CBM), high-level human-interpretable concepts are first extracted through an encoder $g$ and then used to predict the target. Although CBMs are semantically transparent, the causal structure of the model’s inference assumes a straightforward causal structure where concepts are causally independent and are all direct causes of the target. (c) In Causal Concept Graph Models (Causal CGMs), both the concepts’ semantics and the inference’s causal structure are transparent.
Although CBMs have been shown to be semantically transparent (Kim et al., 2018; Facchini & Termine, 2021) [^1] Here by semantic transparent we mean a model that focuses on features that are meaningful and understandable by human experts in the model’s application domain., in this work we argue that these models still rely on two unrealistic assumptions (visualised in Fig. 1(b)). First, they assume that concepts are causally independent, an assumption that breaks in most real-world tasks. For instance, CBMs might treat “smoking” and other health conditions as independent predictors of an insurance premium, ignoring how intervening on smoking habits might have a significant impact on health conditions. Second, and perhaps more importantly, CBMs operate with the underlying assumption that concepts are the direct causes of the model’s predictions. These assumptions oversimplify and constrain the model’s inference structure, preventing the model from using more complex and effective concept dependencies (Beckers, 2022). As a result, the problem of causal opacity still represents a key open challenge at the intersection of DL, causality, and XAI.
Contributions
@@
In response to this challenge, in this paper we introduce Causal Concept Graph Models (Causal CGMs, see Figure 1(c)). Causal CGMs are a new a concept-based architecture whose decision-making process is causally transparent by design, avoiding the unrealistic assumption that concepts must be causally independent and direct causes of the class prediction. The results of our experiments show that Causal CGMs can: (i) match the generalisation performance of state-of-the-art causally-opaque models, (ii) enable human-in-the-loop corrections of mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also the accuracy of the explanation provided for a specific instance, and (iii) support the analysis of both interventional and counterfactual questions, thereby improving the model’s causal interpretability and supporting the effective verification of its reliability and fairness.
@@ 2 Background
A causal model (CM) is a mathematical representation of the mechanisms (rules, laws) that explain how different variables of a given target-system causally influence each other. In computer science and statistics, the standard framework for causal modelling is that of Structural Causal Models (SCM) proposed by Pearl (2009). Formally, an SCM $\mathcal{M}$ is a triplet $(\mathcal{U},\mathcal{V},\mathcal{F}_{\theta})$ where: $\mathcal{U}$ is a set of exogenous variables representing latent factors with a causal influence on the modelled target-system; $\mathcal{V}$ is a set of endogenous variables representing observable and measurable variables; $\mathcal{F}_{\theta}$ is a set of functions (or structural equations) parametrized by $\theta$ describing the causal mechanisms, which determine the values of each endogenous variable $v_{i}\in\mathcal{V}$ by computing the conditional probability $p(v_{i}|u_{i},\text{pa}(v_{i});\theta)$, where $u_{i}\in\mathcal{U}$ and $\text{pa}(v_{i})$ are, respectively, the set of exogenous and endogenous variables causally affecting $v_{i}\in\mathcal{V}$. Every SCM can be associated with a directed acyclic graph (DAG) whose nodes represent variables and edges represent direct causal connections. Interventions can be modelled in SCMs through the do-operator (Pearl, 1995; 2009) . This operator induces a modification on the model’s structure by changing the value of an endogenous variable which is fixed to a value $\kappa\in\mathbb{R}$ by external means, effectively breaking its original causal dependencies within the model. Formally, applying $do(v_{i}=\kappa)$ to a model $\mathcal{M}$ results in a new model $\mathcal{M}_{v_{i}=\kappa}$ which is identical to $\mathcal{M}$ except that the equation for $v_{i}$ is replaced with a constant value $\kappa$, removing all arrows into $v_{i}$ in the model annexed DAG.
@@
Concept-based models (Kim et al., 2018; Chen et al., 2020; Yeh et al., 2020; Kim et al., 2023; Barbiero et al., 2023; Oikarinen et al., 2023) are interpretable architectures that explain their predictions using high-level units of information (i.e., “concepts”). Most of these approaches can be formulated as a Concept Bottleneck Model (CBM) (Koh et al., 2020). Given a sample’s raw features $\mathbf{x}\in X\subseteq\mathbb{R}^{d}$ (e.g., an image’s pixels), a set of $r$ concepts $c_{i}\in C\subseteq\{0,1\}^{r}$ (e.g., “red”, “round”), and a set of $l$ class labels $y_{j}\in Y\subseteq\{0,1\}^{l}$ (e.g., labels “apple” or “tomato”), a CBM estimates the conditional distribution $\prod_{j}p(y_{j}\mid\text{pa}(y_{j})=c_{1},\dots,c_{r};\theta_{f})\prod_{i}p(c_{i}\mid\mathbf{x};\theta_{g})$ where $p(c_{i}\mid\mathbf{x};\theta_{g})$ is a set of independent Bernoulli distributions. This formulation emphasises two key CBMs’ assumptions: i) concepts are direct causes of class predictions, and ii) concepts are conditionally independent of one another. Usually, concept-based models represent a concept $c_{i}$ using its predicted probability $\hat{c}_{i}\in[0,1]$. This representation, however, might significantly degrade task accuracy when the provided concepts are incomplete (Mahinpei et al., 2021). To overcome this issue, Concept Embedding Models (CEMs) (Espinosa Zarlenga et al., 2022) use high-dimensional embeddings $\mathbf{\hat{c}}_{i}\in\mathbb{R}^{z}$ to represent concepts alongside their truth degrees $\hat{c}_{i}\in[0,1]$.
3 Causal Concept Graph Models
Problem: This work addresses causal opacity denotes the difficulty in understanding the hidden causal structure behind a DL model’s inference and decision-making process. Research question: To overcome this, we aim to design DL models where each prediction can be traced back to a chain of semantically meaningful causes. Challenged assumptions: In pursuing this goal, we challenge two key assumptions of Concept Bottleneck Models (CBMs): i) humans know a priori which interpretable variables (i.e., concepts) are direct causes of a given class label (and the dataset embeds this prior knowledge), and ii) all interpretable variables used for explanations (i.e., concepts) are conditionally independent. Solution: We introduce Causal Concept Graph Models (Causal CGMs), a class of models making class predictions through cause-effect chains of interpretable variables. In what follows, we first introduce Causal CGMs and then proceed to show how their decision-making process is causally transparent. Finally, we conclude this section by describing their learning process.
@@ 3.1 Blueprint
Dropping the assumptions of CBMs—that concepts are independent and direct causes of class labels—requires a model able to capture all possible dependencies [^2] This naive formulation would lead to a cyclic probabilistic model, which becomes computationally intractable due to the difficulty in handling cycles in probabilistic models. Causal Concept Graph Models tackle this issue by ensuring the learning process is efficient and produces simple explanations by design (as described in Sec. 3.3). between interpretable variables (i.e., any classification label). To model such dependencies, Concept Graph Models duplicate the layer of interpretable variables $\mathcal{V}$ by introducing an additional layer of identical copies $\mathcal{V}^{\prime}$ [^3] A detailed analysis of this transformation, including the “unfolding” of the original cyclic model, is provided in App. B and a neural parametrization in App. D.. Using this additional layer, Causal CGMs make predictions of each $v_{i}\in\mathcal{V}$ using as possible inputs all $v_{j}^{\prime}\in\mathcal{V}^{\prime}$ for all $j\neq i$. This inference structure enables CGMs to trace each prediction in $\mathcal{V}$ back to a set of interpretable variables in $\mathcal{V}^{\prime}$. As a result, the set of copies $\mathcal{V}^{\prime}$ in a Causal CGM play the role of “explaining variables”, akin to a CBM’s concepts, while the set of variables $\mathcal{V}$ play the role of “explained variables”, akin to a CBM’s tasks.
Definition 3.1 (Causal Concept Graph Model).
Given an observed feature vector $\mathbf{x}$, a set of $k\in\mathbb{N}$ latent factors $u_{i}\in\mathcal{U}$ each associated with a pair of identical high-level interpretable variables $v_{i}\in\mathcal{V}$ and $v_{i}^{\prime}\in\mathcal{V}^{\prime}$ with $v_{i}=v_{i}^{\prime},\ \forall i\in\{1,\dots,k\}$, a Causal Concept Graph Model $\Gamma=(\mathcal{N},\mathcal{E},\mathcal{F}_{\theta})$ with
  • nodes $\mathcal{N}=\{x\}\cup\mathcal{U}\cup\mathcal{V}^{\prime}\cup\mathcal{V}$,
  • edges =$\mathcal{E}=\big{\{}(\mathbf{x},u_{i})\mid u_{i}\in\mathcal{U}\big{\}}\cup\big{\{}(u_{i},v_{i}^{\prime})\mid i\in\{1,\cdots,k\}\big{\}}\cup\big{\{}(u_{i},v_{i})\mid i\in\{1,\cdots,k\}\big{\}}\cup\{(v_{i}^{\prime},v_{j})\mid v_{i}^{\prime}\in\mathcal{V}^{\prime},v_{j}\in\mathcal{V},i\neq j\}$,
  • and mechanisms $\mathcal{F}_{\theta}=\{\zeta,s,f\}$,
represents the joint conditional distribution:
${p(v_{1}\dots,v_{k},v_{1}^{\prime},\dots,v_{k}^{\prime},u_{1},\dots,u_{k}\mid x;\theta)=\prod_{i=1}^{k}\overbrace{p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i};\theta_{f})}^{\text{endogenous predictor}}\prod_{j\neq i}\overbrace{p(v_{j}^{\prime}\mid u_{j};\theta_{s})}^{\text{copies predictor}}\overbrace{p(u_{j}\mid\mathbf{x};\theta_{\zeta})}^{\text{exogenous encoder}}}$
@@
3.2 Causal transparency in Concept Graph Models
Notice how the distribution $p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i};\theta_{f})$ can be associated with a structural causal model $\mathcal{M}_{CGM}=(\{u_{i}\},\mathcal{V}^{\prime}\cup\{v_{i}\},\{f_{i}\})$ with causal mechanism $f_{i}$ where $u_{i}\in\mathcal{U}$ is an exogenous variable representing latent, uninterpretable information (e.g., noise), $v_{i}\in\mathcal{V}$ is an endogenous variable representing interpretable, symbolic information, and $f_{i}:U\times V\to V$ is a function describing the causal mechanism that determines the value of $v_{i}$ given its parents $\text{pa}_{\mathcal{V}^{\prime}}(v_{i})$. Such dependencies are captured by a graph $\mathcal{G}^{\prime}=(\mathcal{V}\cup\mathcal{V}^{\prime},\{(v_{i}^{\prime},v_{j})\mid v_{i}^{\prime}\in\mathcal{V}^{\prime},v_{j}\in\mathcal{V},v_{i}^{\prime}\neq v_{j}\})$, representing all direct causal dependencies between endogenous variables. By modelling such dependencies, and considering that by Def. 3.1 the variables $v_{i}$ and $v_{i}^{\prime}$ are identical copies of the same interpretable variable, Causal CGMs allow any endogenous variable $v_{i}$ to be expressed in terms of its parents in $\mathcal{V}$, denoted as $\text{pa}_{\mathcal{V}}(v_{i})$, rather than in terms of its parents in $\mathcal{V}^{\prime}$, i.e., $\text{pa}_{\mathcal{V}^{\prime}}(v_{i})$, as formalised in the following theorem (proof in App. C).
Theorem 3.2.
Given a Causal Concept Graph Model $\Gamma=(\mathcal{N},\mathcal{E},\mathcal{F}_{\theta})$, let the set of endogenous root nodes be defined as: $\text{roots}(\mathcal{G}^{\prime})=\{v_{i}^{\prime}\in\mathcal{V}^{\prime}\mid\nexists(v_{j}^{\prime},v_{i})\in\mathcal{E}_{\mathcal{G}^{\prime}}\}$, and the set of children of root nodes as: $\text{ch}(\text{roots}(\mathcal{G}^{\prime}))=\{v_{i}\in\mathcal{V}\mid\exists(v_{j}^{\prime},v_{i})\in\mathcal{E}_{\mathcal{G}^{\prime}},v_{j}^{\prime}\in\text{roots}(\mathcal{G}^{\prime})\}$. Then, for all nodes in the set $\mathcal{V}\setminus\{\text{roots}(\mathcal{G}^{\prime})\cup\text{ch}(\text{roots}(\mathcal{G}^{\prime}))\}$, i.e., for all nodes that are neither root nodes nor children of root nodes, the following holds:
$p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i};\theta_{f})=p(v_{i}\mid\text{pa}_{\mathcal{V}}(v_{i}),u_{i};\theta_{f})$